Circular No. 5965

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

OCCULTATION BY (2060) CHIRON J. L. Elliot, Massachusetts Institute of Technology (MIT) and Lowell

Observatory; E. W. Dunham, Ames Research Center, NASA; and C. B. Olkin, MIT, report on the prediction and successful Kuiper Airborne Observatory (KAO) observation of the Mar. 9 occultation of the 11.9-mag star Ch08 (Bus et al., A.J., in press) by (2060) Chiron: “CCD-strip-scan observations were made by C. Ford (SETI Institute) and R. P. S. Stone (Lick Observatory) using the Crossley telescope at Lick. These data were reduced by S. W. McDonald, R. M. Bandyopadhyay, and their colleagues at MIT. The KAO optical light curve shows one integration interval of 0.5 s (at 23^{h}28^{m}55^{s} UT, when the KAO was near Recife) to have a drop of ∼60 percent, with lesser drops in several neighboring intervals on either side. This is similar in character to the occultation light curve for Ch02 obtained from Palomar on 1993 Nov. 7 (IAUC 5898). Simultaneous K-band observations were made but have not been reduced yet. The occulting object was almost certainly some structure within the coma near the nucleus rather than the nucleus itself. Other members of the KAO observing team were D. K. Gilmore, D. M. Rank, and P. Temi (University of California at Santa Cruz and Lick Observatory); D. Lazzaro (Observatorio Nacional) participated as Brazil’s primary scientific representative.”

W. B. Hubbard, University of Arizona, writes: “Ground-based observations of the Mar. 9 occultation by Chiron were attempted in Brazil at the Pico dos Dias Observatory, Itajubá, by H. Reitsema (Ball Aerospace), A. Barucci (Paris Observatory), and J. Barroso (Observatorio Nacional), and by three mobile teams from MIT, Lowell Observatory, and the University of Arizona, working in collaboration with the Observatorio Nacional and the Paris Observatory (S. J. Bus, D. F. Lopes, M. Buie, B. Sicardy, R. Marcialis, D. W. Foryta), but no data were gathered due to clouds. At the South African Astronomical Observatory, Sutherland, conditions were photometric, and D. Kurtz (University of Cape Town) observed the occultation in Johnson V with the 0.5-m telescope, recording a single brief and deep event at ∼ 23^{h}23^{m}45^{s} UTC. The event lasted <0.5 s, with a maximum stellar signal drop of ∼75 percent; it resembles the unresolved dips detected during the 1993 Nov. 7 Chiron occultation (IAUC 5898) and may be associated with an opaque feature a few km wide but extending some hundreds of km from the nucleus.”

1994 April 5 Daniel W. E. Green